OpenStudy (anonymous):

Find the equation, in standard form, of the line passing through the points (2,-3) and (4,2). a. 2x-3y=9 b. 5x+2y=16 c. 5x-2y=16

OpenStudy (anonymous):

first find the slope of the line: 2+3 5 -------- = -------- 4-2 2 Now you have : y = 5/2x + b. substitute one of the points into the line, and you get: 2 = 5/2 * 4 + b 2= 10 + b b= -8 this would make our equation: y = 5/2x -8. because this doesn't equal any of the equations in the list, we need to convert them so that they're in point slope form by isolating the y variable. I will omit this step as I can see that the answer is letter A.